Slope of 6y-2x=36?
Answer by nerdybill(7384)

Slope of 6y-2x=36?
.
Rearrange so that it is in "slope-intercept" form
y = mx + b
.
6y-2x=36
6y = 2x+36
dividing both sides by 6:
y = (1/3)x + 6
.
Slope is 1/3